# GhostEI-Bench: Do Mobile Agents Resilience to Environmental Injection in Dynamic On-Device Environments?

### v2

@@ Abstract
Vision-Language Models (VLMs) are increasingly deployed as autonomous agents to navigate mobile Graphical User Interfaces (GUIs). However, their operation within dynamic on-device ecosystems, which include notifications, pop-ups, and inter-app interactions, exposes them to a unique and underexplored threat vector: environmental injection. Unlike traditional prompt-based attacks that manipulate textual instructions, environmental injection contaminates the agent’s visual perception by inserting adversarial UI elements, such as deceptive overlays or spoofed notifications, directly into the GUI. This bypasses textual safeguards and can derail agent execution, leading to privacy leakage, financial loss, or irreversible device compromise. To systematically evaluate this threat, we introduce GhostEI-Bench, the first benchmark dedicated to assessing mobile agents under environmental injection attacks within dynamic, executable environments. Moving beyond static image-based assessments, our benchmark injects adversarial events into realistic application workflows inside fully operational Android emulators, assessing agent performance across a range of critical risk scenarios. We also introduce a novel evaluation protocol where a judge LLM performs fine-grained failure analysis by reviewing the agent’s action trajectory alongside the corresponding sequence of screenshots. This protocol identifies the precise point of failure, whether in perception, recognition, or reasoning. Our comprehensive evaluation of state-of-the-art agents reveals their profound vulnerability to deceptive environmental cues. The results demonstrate that current models systematically fail to perceive and reason about manipulated UIs. GhostEI-Bench provides an essential framework for quantifying and mitigating this emerging threat, paving the way for the development of more robust and secure embodied agents.
@@ 1 Introduction
The growing autonomy of these agents inevitably brings their security and trustworthiness into sharp focus. In response, a foundational body of research has emerged to evaluate their robustness. Dedicated efforts on mobile platforms, such as MobileSafetyBench (Lee et al., 2024), alongside broader frameworks like MLA-Trust (Yang et al., 2025b) and MMBench-GUI (Wang et al., 2025b), have provided critical insights into agent reliability and policy adherence. However, these pioneering evaluations primarily assess an agent’s response to predefined, static threats. Their methodologies, often focused on analyzing static UI states or refusing harmful textual prompts, possess a critical blind spot for hazards that emerge dynamically and unpredictably.
@@
However, these existing evaluations largely overlook a more insidious threat vector unique to the interactive and unpredictable nature of mobile ecosystems: dynamic environmental injection. This attack surface is fundamentally different from previously studied risks. Here, adversaries inject unexpected, misleading UI elements—deceptive pop-ups, spoofed notifications, or malicious overlays—directly into the environment during an agent’s task execution. Such attacks exploit the agent’s primary reliance on visual perception, bypassing textual safeguards entirely to corrupt its decision-making process at a critical moment. While nascent studies have begun to demonstrate the feasibility of these attacks (Zhang et al., 2024; Yang et al., 2025a; Liu et al., 2025b), they often lack a unified framework for systematic evaluation. This critical gap leaves the resilience of mobile agents against real-time, unpredictable interruptions dangerously unquantified, posing severe risks of privacy leakage, financial fraud, and irreversible device compromise.
@@
To address this critical gap, we introduce GhostEI-Bench, the first comprehensive benchmark designed to systematically evaluate mobile agent robustness against dynamic environmental injection attacks in fully operational on-device environments(see 1 for an overview). Moving beyond static, image-based assessments, GhostEI-Bench injects a wide range of adversarial events into realistic, multi-step task workflows within Android emulators. These scenarios, spanning seven critical risk fields and diverse application domains, are designed to manipulate agent perception and action at runtime. We further propose a novel evaluation protocol where a judge LLM performs fine-grained failure analysis on the agent’s action trajectory. GhostEI-Bench provides an essential framework for quantifying and mitigating this emerging threat, paving the way for the development of more robust and secure embodied agents that can be safely deployed in the real world. Drawing from the GhostEI-Bench framework, we conduct a comprehensive evaluation of 8 prominent VLM agents against environmental injection attacks, uncovering severe security vulnerabilities. We find that for many models, the Vulnerability Rate falls within the 40% to 55% range. Beyond this baseline evaluation, we also explore how incorporating self-reflection and reasoning modules affects agent robustness.
Overall, this work makes the following contributions:
@@ 2.2 Adversarial Vulnerabilities in Multimodal Agents
Recent studies reveal that multimodal and LLM-based agents remain broadly vulnerable to adversarial manipulations across both visual and semantic channels. Wu et al. (2024) show that even minor visual perturbations can reliably misguide multimodal agents’ decision pathways. Aichberger et al. (2025) demonstrate that small, human-imperceptible OS-level patches can hijack GUI agents without affecting interface usability. Beyond visual cues, Wang et al. (2025a) introduce cross-modal prompt injections where crafted images override text-based reasoning and safety constraints. In the language domain, Fu et al. (2024) present Imprompter, revealing that structured linguistic sequences can coerce agents into unsafe or unintended tool operations. Cheng et al. (2025a) identify a training-stage backdoor mechanism in which agents are conditioned to trigger harmful actions under particular historical states, highlighting that adversarial risks extend beyond inference into the full development pipeline. These works collectively underscore the inherent fragility of the underlying models driving GUI agents.
@@ 2.3 Environmental Injection Attacks
Within the specific context of GUI agents, environmental dynamics serve as one viable pathway for adversarial risks. The environmental dynamics is prevalent in GUI agents on both smartphones and computers, where notifications, pop-ups, and advertisements continually interfere decision contexts. Ma et al. (2025) highlight that multimodal agents are inherently susceptible to such environmental distractions, where visual noise can easily derail the reasoning process. In this vein, Zhang et al. (2024) demonstrate that adversarial pop-ups can significantly disrupt the task execution of VLM agents in environments such as OSWorld (Xie et al., 2024) and VisualWebArena (Koh et al., 2024), effectively circumventing common defensive prompts and revealing critical vulnerabilities in agent architectures. RiOSWorld (Yang et al., 2025a) categorizes such phenomena as environmental risks and introduces evaluators for intention and completion under dynamic threats. AgentHazard (Liu et al., 2025b) further shows that scalable GUI-hijacking via third-party content can systematically mislead mobile agents. More recently, Chen et al. define the concept of Active Environmental Injection Attacks (AEIA) and demonstrate that adversarial notifications can effectively exploit reasoning gaps in AndroidWorld (Chen et al., 2025). These attacks are often perceptually conspicuous yet strategically deceptive, specifically designed to divert agent behavior through seemingly imperative cues. While these studies often evaluate an agent’s robustness against specific, single injection risks in isolation, they have firmly established the feasibility and severity of dynamic injection attacks, highlighting the significant challenge they pose for comprehensively evaluating the safety of GUI agents.
@@ 3.1 Design Principles and Threat Model
GhostEI-Bench is designed to provide a comprehensive and rigorous evaluation of mobile agent safety. Its core design leverages executable Android emulators to create interactive and realistic environments, with snapshot-based mechanisms to ensure reproducibility.See Figure  2 for an overview of the unified threat model and the three attack vectors evaluated in our benchmark.
@@ 3.2 Benchmark Construction and Composition
This principled approach ensures that our benchmark is comprehensive, based on the cross product of three core dimensions: representative domains, critical risk fields, and specific attack vectors.To systematically populate this design space, we first utilized a Large Language Model (LLM) to programmatically generate a diverse set of test cases, with each case representing a unique permutation of these dimensions and conforming to the JSON structure defined in Appendix  B.2. Each auto-generated task then underwent a rigorous manual review by human experts. This curation process involved verifying the scenario’s feasibility and realism within the target application, ensuring the logical coherence between the prompt, content, and result fields, and confirming the accurate alignment of the critical risk fields. For dynamic attacks, experts paid special attention to enforcing the decoupling between the benign user prompt and the malicious payload. Upon a task’s validation, the same experts proceeded to configure the necessary environmental preconditions, such as creating and placing files as specified in the files name field, to ensure the task was fully executable.
@@ B.2 Data Format Specification
A complete example of a data entry corresponding to an "Overlay Attack" is shown in Listing 1.
@@ Appendix E Detailed Prompt
This section presents the detailed prompt provided to the Judge Language Model. This model is tasked with evaluating the agent’s behavior for security vulnerabilities and task completion based on the provided inputs and criteria. The complete prompt is shown in Listing 2.

### v3

@@ Abstract
Vision-Language Models (VLMs) are increasingly deployed as autonomous agents to navigate mobile Graphical User Interfaces (GUIs). However, their operation within dynamic on-device ecosystems, including notifications, pop-ups, and inter-app interactions, exposes them to a unique and underexplored threat vector: environmental injection. Unlike traditional prompt-based attacks that manipulate textual instructions, environmental injection contaminates the agent’s visual perception by inserting adversarial UI elements, such as deceptive overlays or spoofed notifications, directly into the GUI. This bypasses textual safeguards and can derail agent execution, leading to privacy leakage, financial loss, or irreversible device compromise. To systematically evaluate this threat, we introduce GhostEI-Bench, the first benchmark dedicated to assessing mobile agents under environmental injection attacks within dynamic, executable environments. Moving beyond static image-based assessments, our benchmark injects adversarial events into realistic application workflows inside fully operational Android emulators, assessing agent performance across a range of critical risk scenarios. We also introduce a novel evaluation protocol where a judge LLM performs fine-grained failure analysis by reviewing the agent’s action trajectory alongside the corresponding sequence of screenshots. This protocol identifies the precise point of failure, whether in perception, recognition, or reasoning. Our comprehensive evaluation of state-of-the-art agents reveals their profound vulnerability to deceptive environmental cues. The results demonstrate that current models systematically fail to perceive and reason about manipulated UIs. GhostEI-Bench provides an essential framework for quantifying and mitigating this emerging threat, paving the way for the development of more robust and secure embodied agents. The project is available at https://github.com/cyChen2003/Ghost-EI.
@@ 1 Introduction
The growing autonomy of these agents inevitably brings their security and trustworthiness into sharp focus (Liu et al., 2026). In response, a foundational body of research has emerged to evaluate their robustness. Dedicated efforts on mobile platforms, such as MobileSafetyBench (Lee et al., 2024), alongside broader frameworks like MLA-Trust (Yang et al., 2025b) and MMBench-GUI (Wang et al., 2025b), have provided critical insights into agent reliability and policy adherence. However, these pioneering evaluations primarily assess an agent’s response to predefined, static threats. Their methodologies, often focused on analyzing static UI states or refusing harmful textual prompts, possess a critical blind spot for hazards that emerge dynamically and unpredictably.
@@
However, these existing evaluations largely overlook a more insidious threat vector unique to the interactive and unpredictable nature of mobile ecosystems: dynamic environmental injection. This attack surface is fundamentally different from previously studied risks. Here, adversaries inject unexpected, misleading UI elements—deceptive pop-ups, spoofed notifications, or malicious overlays—directly into the environment during an agent’s task execution. Such attacks exploit the agent’s primary reliance on visual perception, bypassing textual safeguards entirely to corrupt its decision-making process at a critical moment. While nascent studies have begun to demonstrate the feasibility of these attacks (Zhang et al., 2024; Yang et al., 2025a; Liu et al., 2025b; Li et al., 2025), they often lack a unified framework for systematic evaluation. This critical gap leaves the resilience of mobile agents against real-time, unpredictable interruptions dangerously unquantified, posing severe risks of privacy leakage, financial fraud, and irreversible device compromise.
@@
To address this critical gap, we introduce GhostEI-Bench, the first comprehensive benchmark designed to systematically evaluate mobile agent robustness against dynamic environmental injection attacks in fully operational on-device environments (see Figure 1 for an overview). Moving beyond static, image-based assessments, GhostEI-Bench injects a wide range of adversarial events into realistic, multi-step task workflows within Android emulators. These scenarios, spanning seven critical risk fields and diverse application domains, are designed to manipulate agent perception and action at runtime. We further propose a novel evaluation protocol where a judge LLM performs fine-grained failure analysis on the agent’s action trajectory. GhostEI-Bench provides an essential framework for quantifying and mitigating this emerging threat, paving the way for the development of more robust and secure embodied agents that can be safely deployed in the real world. Drawing from the GhostEI-Bench framework, we conduct a comprehensive evaluation of 8 prominent VLM agents against environmental injection attacks, uncovering severe security vulnerabilities. We find that for many models, the Vulnerability Rate falls within the 40% to 55% range. Beyond this baseline evaluation, we also explore how incorporating self-reflection and reasoning modules affects agent robustness. Overall, this work makes the following contributions:
@@ 2.2 Adversarial Vulnerabilities in Multimodal Agents
Recent studies reveal that multimodal and LLM-based agents remain broadly vulnerable to adversarial manipulations across both visual and semantic channels. Wu et al. (2024) show that even minor visual perturbations can reliably misguide multimodal agents’ decision pathways. Aichberger et al. (2025) demonstrate that small, human-imperceptible OS-level patches can hijack GUI agents without affecting interface usability. Beyond visual cues, Wang et al. (2025a) introduce cross-modal prompt injections where crafted images override text-based reasoning and safety constraints. In the language domain, Fu et al. (2024) presents Imprompter, revealing that structured linguistic sequences can force agents into unsafe or unintended tool operations. Cheng et al. (2025a) identify a training-stage backdoor mechanism in which agents are conditioned to trigger harmful actions under particular historical states, highlighting that adversarial risks extend beyond inference into the full development pipeline. These works collectively underscore the inherent fragility of the underlying models driving GUI agents.
@@ 2.3 Environmental Injection Attacks
Within the specific context of GUI agents, environmental dynamics serve as one viable pathway for adversarial risks. The environmental dynamics is prevalent in GUI agents on both smartphones and computers, where notifications, pop-ups, and advertisements continually interfere decision contexts. Ma et al. (2025) highlight that multimodal agents are inherently susceptible to such environmental distractions, where visual noise can easily derail the reasoning process. In this vein, Zhang et al. (2024) demonstrate that adversarial pop-ups can significantly disrupt the task execution of VLM agents in environments such as OSWorld (Xie et al., 2024) and VisualWebArena (Koh et al., 2024), effectively circumventing common defensive prompts and revealing critical vulnerabilities in agent architectures. RiOSWorld (Yang et al., 2025a) categorizes such phenomena as environmental risks and introduces evaluators for intention and completion under dynamic threats. AgentHazard (Liu et al., 2025b) further shows that scalable GUI-hijacking via third-party content can systematically mislead mobile agents. More recently, Chen et al. (2025) define the concept of Active Environmental Injection Attacks (AEIA) and demonstrate that adversarial notifications can effectively exploit reasoning gaps in AndroidWorld. These attacks are often perceptually conspicuous yet strategically deceptive, specifically designed to divert agent behavior through seemingly imperative cues. While these studies often evaluate an agent’s robustness against specific single injection risks in isolation, they have firmly established the feasibility and severity of dynamic injection attacks, highlighting the significant challenge they pose for comprehensively evaluating the safety of GUI agents.
@@ 3.1 Design Principles and Threat Model
GhostEI-Bench is designed to provide a comprehensive and rigorous evaluation of mobile agent safety. Its core design leverages executable Android emulators to create interactive and realistic environments, with snapshot-based mechanisms to ensure reproducibility. See Figure 2 for an overview of the unified threat model and the three attack vectors evaluated in our benchmark.
@@ 3.2 Benchmark Construction and Composition
This principled approach ensures that our benchmark is comprehensive, based on the cross product of three core dimensions: representative domains, critical risk fields, and specific attack vectors.To systematically populate this design space, we first utilized a Large Language Model (LLM) to programmatically generate a diverse set of test cases, with each case representing a unique permutation of these dimensions and conforming to the JSON structure defined in Appendix B.2. Each auto-generated task then underwent a rigorous manual review by human experts. This curation process involved verifying the scenario’s feasibility and realism within the target application, ensuring the logical coherence between the prompt, content, and result fields, and confirming the accurate alignment of the critical risk fields. For dynamic attacks, experts paid special attention to enforcing the decoupling between the benign user prompt and the malicious payload. Upon a task’s validation, the same experts proceeded to configure the necessary environmental preconditions, such as creating and placing files as specified in the files name field, to ensure the task was fully executable.
@@ 5 Conclusion
6 ETHICS STATEMENT
This paper evaluates mobile GUI agents under synthetic environmental injections executed exclusively in controlled Android emulators. No human subjects or real personal data were involved; credentials, contacts, and content are fictitious, and all ‘send/submit’ controls are inert placeholders, no scripts and no network side effects. While our analysis may reveal failure modes that could be misused, our intent is defensive. We follow provider terms for any third-party models and use outputs only for aggregate reporting. The work aims to help the community measure and reduce risk in agentic systems.
Acknowledgments
This work was sponsored by the Joint Funds of the National Natural Science Foundation of China (Grant No.U21B2020) , the China Postdoctoral Science Foundation (Grant No.2025M771514), and Shanghai Artificial Intelligence Laboratory.
Appendix
@@ B.2 Data Format Specification
A complete example of a data entry corresponding to an “Overlay Attack” is shown in 1.
@@ Appendix E Detailed Prompt
This section presents the detailed prompt provided to the Judge Language Model. This model is tasked with evaluating the agent’s behavior for security vulnerabilities and task completion based on the provided inputs and criteria. The complete prompt is shown in 2.